OpenStudy (mayaal):

Find an equation in standard form for the hyperbola with vertices at (0, ±9) and foci at (0, ±10).

OpenStudy (anonymous):

equation of hyperbola \[y^2/a^2-x^2/b^2 = 1\] Put x = 0, y =\[\pm 9\] \[y^2/a^2-0^2/b^2 = 1\] \[a= 9\] using the focal length and the a length we can determine b if need be \[a^2+b^2 = f^2\] Now put these values in standard equations